### v1

@@ 2.3 MinHashLSH for Scalable MinHash
Figure 1: Breakdown of wall clock time on 5% of peS2o for a conventional LSH algorithm and our LSHBloom method
@@ 4.1 Overview
We propose a memory-efficient alternative to MinHashLSH for document deduplication. In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for hash collisions. The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of Bender et al. (2022).
@@ 4.2 Insertion
When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers onto a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value is computed using a universal hash function, with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values Carter & Wegman (1977). Then, we can hash a vector $\bar{x}$ of $r$ integers as follows:
@@ 4.4 Advantages
While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration. The additional false positive probability scales with the number of documents we intend to insert and scales inversely with the size of each filter. Therefore, there is a trade-off between memory usage and correctness that is mediated by the size of each filter. However, even for very large datasets of, say, 100M documents, this does not prove to be an issue as we can set an effective false positive probability overhead as low as 1e-15 and only incur a space cost of 8.5 GB for our index compared to 460 GB for the equivalent LSHIndex.
@@ 5.2.1 LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figures 2(a) and 2(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 25K dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate. However, there is a clear saturation point at 128 permutation functions, after which we observe only marginal performance increases. This is particularly relevant for LSHBloom, for which increasing the number of permutation functions blindly can lead to slight degradations in performance—due to the introduction of more bands in the signature matrix by the LSH function, requiring more Bloom filters, and therefore leading to an increase in the false positive rate. This can be avoided, however, by jointly tuning the false positive rate and the number of permutation functions. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.6 yields the best performance, as more stringent thresholds lead to a diminishing F1 score.
@@ 5.2.2 Paragraph Level and N-Gram Methods
Figure 4 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score. Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. N-gram methods such as DOLMA-ngram and CCNet do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques. Figures 3(a) and 3 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is similar to the notion of evaluating the set overlap between documents but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, both paragraph and n-gram methods demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates. In contrast, n-gram methods yield many false negatives at high thresholds and yield many false positives at low thresholds.
@@ 5.3.2 Estimated Resource Usage at Scale
Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance. Figures 5(a) and 5(b) plot the maximum F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. Additionally, LSHBloom displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown in Section 5.3.1, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.

### v2

@@ 2.3 MinHashLSH for Scalable MinHash
Figure 1: Breakdown of wall clock time on 5% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
@@ 4.1 Overview
We propose a memory-efficient alternative to MinHashLSH for document deduplication. In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates. The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of Bender et al. (2022).
@@ 4.2 Insertion
When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value is computed using a universal hash function (in our case, SHA1 3rd & Jones (2001)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values. Then, we can hash a vector $\bar{x}$ of $r$ integers as follows Carter & Wegman (1977):
@@ 4.3 Querying
4.4 Error Rate
In traditional MinHashLSH, a collision in any band causes a document to be marked as a candidate duplicate. This is also the case in LSHBloom. However, we make use of Bloom filters, which incur additional false positive overhead. Here, we analyze the error rate of LSHBloom. A false positive is a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom, and a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
[…]
The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows. Leskovec et al. (2014) determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
$FP=\int_{0}^{T}1-(1-t^{r})^{b}dt$ (1)
$FN=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$ (2)
Zhu et al. (2016) demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
A Bloom filter will not yield false negatives, so the overall false negative rate is unaffected in LSHBloom. If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result. If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$. Therefore, the overall false positive rate for LSHBloom is given by:
$FP_{total}=FP+(1-FP)*p_{effective}$
Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible.
@@ 4.5 Advantages
While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration. This is another advantage of our technique: we can observe massive gains in space while only introducing a single additional hyperparameter, meaning that our hyperparameter tuning burden is comparable to that of MinHashLSH.
The additional false positive error overhead scales with the number of documents we intend to insert and scales inversely with the size of each filter. Therefore, there is a trade-off between memory usage and correctness that is mediated by the size of each filter. However, even for very large datasets of, say, 100M documents, this does not prove to be an issue as we can set an effective false positive probability overhead as low as 1e-15 and only incur a space cost of 8.5 GB for our index compared to 460 GB for the equivalent LSHIndex.
@@ 5.1.5 Choice of values for Tuning Parameters
Figure 2: Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given $N$=24,956 documents and $p=$1e-5
@@ 5.2.1 LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 25K dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate. However, there is a clear saturation point at 128 permutation functions, after which we observe only marginal performance increases. This is particularly relevant for LSHBloom, for which increasing the number of permutation functions blindly can lead to slight degradations in performance—due to the introduction of more bands in the signature matrix by the LSH function, requiring more Bloom filters, and therefore leading to an increase in the false positive rate. This can be avoided, however, by jointly tuning the false positive rate and the number of permutation functions. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.6 yields the best performance, as more stringent thresholds lead to a diminishing F1 score.
@@ 5.2.2 Paragraph Level and N-Gram Methods
Figure 5 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score. Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. N-gram methods such as DOLMA-ngram and CCNet do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques. Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is similar to the notion of evaluating the set overlap between documents but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, both paragraph and n-gram methods demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates. In contrast, n-gram methods yield many false negatives at high thresholds and yield many false positives at low thresholds.
@@ 5.3.2 Estimated Resource Usage at Scale
Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance. Figures 6(a) and 6(b) plot the maximum F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. Additionally, LSHBloom displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown in Section 5.3.1, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.